# Knowledge Adaptation as Posterior Correction

### v1

@@ 3.2 Influence Estimation as Posterior Correction
$\boldsymbol{\lambda}_{t\backslash j}\approx\arg\min_{\boldsymbol{\mu}}\,\,-\langle\boldsymbol{\mu},\widetilde{\nabla}\mathbb{E}_{q_{t}}[\ell_{i}]\rangle+\mathbb{D}_{\text{KL}}[q\,\|\,q_{t}]\quad\implies\quad\boldsymbol{\lambda}_{t\backslash j}\approx\boldsymbol{\lambda}_{t}+\widetilde{\nabla}\mathbb{E}_{q_{t}}[\ell_{i}].$
@@
The second expression is obtained by noting that the natural gradient of the KL term is simply $\boldsymbol{\lambda}_{t\backslash j}-\boldsymbol{\lambda}_{t}$ (Khan and Rue, 2023, Eq. 23). This is precisely the Memory-Perturbation Equation shown in Nickl et al. (2023, Eq. 6). Overall, these connections suggests two ways to improve influence estimation: either we can expand the class of posterior family or we could explicitly try to reduce the correction.
@@ 3.3 Model Merging as Posterior Correction
Model merging is a popular technique to improve capabilities of LLMs without retraining them from scratch. Given a pre-trained model with parameters $\boldsymbol{\theta}_{\text{pre}}$ on loss $\ell_{\text{pre}}$, we ‘fine-tune’ models $\boldsymbol{\theta}_{i}$ on individual tasks $\ell_{i}$. For instance, given an English-speaking [^2] Note that this is an anthropomorphization to simplify things. LLMs do not speak like people do. LLM, we may fine-tune it on other languages and merge them all to get a model that can also speak other languages. The merging is done via a simple addition, for instance, Task-Arithmetic (TA) (Ilharco et al., 2023) uses the following merging (with weights $\alpha_{i}>0$):
$\boldsymbol{\theta}_{\text{ta}}=\boldsymbol{\theta}_{\text{pre}}+\sum_{i=1}^{t}\alpha_{i}(\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{\text{pre}}),$ (12)
which, despite its simplicity, works extremely well for reasons unknown. In what follows, we will derive this as a special case of posterior correction with the family $q^{\text{iso}}$. We will then show that better approximation of the correction leads to better merging.
As before, we start by writing merging in a variational formulation denoting the pre-trained model by $q_{\text{pre}}$, for instance, it can be a $q^{\text{iso}}$ candidate with mean $\boldsymbol{\theta}_{\text{pre}}$. Similarly, we denote by $q_{i}$ the model fine-tuned on $\ell_{i}$ but with a prior $q_{\text{pre}}$, that is, by minimizing $\mathbb{E}_{q}[\ell_{i}]+\mathbb{D}_{\text{KL}}[q\,\|\,q_{\text{pre}}]$; we assume both $q_{i}$ to be in the same family as $q_{\text{pre}}$. The use of KL term is similar to the proximal term which is implemented via initialization. The goal is then to obtain the model that is trained from scratch on all the data. A popular strategy for merging posteriors is to use the following rule:
@@
which we term ‘Bayesian Arithmetic (BA)’ to draw an analogy to Task Arithmetic because the two have similar forms with addition and subtraction in TA replaced by multiplication and division in BA. The above rule is commonly used in various distributed Bayesian computation, such as, Bayesian data fusion (Mutambara, 1998; Durrant-Whyte, 2001; Wu et al., 2022), Bayesian Committee Machine (Tresp, 2000), Consensus Monte Carlo (Scott et al., 2022), and approximate inference (Vehtari et al., 2020; Ashman et al., 2022).
Clearly, the posterior merging does not recover the exact posterior fitted jointly, and we will now apply our posterior correction to write an exact expression for the error. We note that $q_{i}$ takes the following dual form for $q_{i}\propto q_{\text{pre}}\exp(-\hat{\ell}_{i|i})$ where the ^$\hat{\ell}_{i|i}$ is the site function for $\ell_{i}$ at $q_{i}$. Similarly, let $q_{\text{pre}}\propto\exp(-\hat{\ell}_{\text{pre}|\text{pre}})$ where $\ell_{\text{pre}}$ is assumed to contain the prior $p_{0}$ used for training from scratch. With this, we can write the retraining of the model jointly on $\ell_{\text{pre}}$, as well as all $\ell_{i}$, as follows,
@@
$\begin{split}q_{\text{retrained}}&=\arg\min_{q\in\mathcal{Q}}\,\,\mathbb{E}_{q}\left[\ell_{\text{pre}}+\sum_{i=1}^{t}\alpha_{i}\ell_{i}\right]+\mathbb{D}_{\text{KL}}[q\,\|\,p_{0}]\\ &=\arg\min_{q\in\mathcal{Q}}\,\,\mathbb{D}_{\text{KL}}[q\,\|\,q_{\text{ba}}]+\sum_{i=1}^{t}\alpha_{i}\underbrace{\mathbb{E}_{q}[\ell_{i}-\hat{\ell}_{i|i}]}_{\text{Task correction}}+\underbrace{\mathbb{E}_{q}[\ell_{\text{pre}}-\hat{\ell}_{\text{pre}|\text{pre}}]}_{\text{LLM correction}}.\end{split}$ (14)
This is obtained in the same way as Eq. 4 but now using the dual forms of $q_{\text{pre}}$ and all $q_{i}$.
It is now easy to check that TA model $\boldsymbol{\theta}_{\text{ta}}$ in Eq. 12 is obtained from $q_{\text{ba}}$ by restricting all posteriors to $q^{\text{iso}}$ family, that is, $q_{i}=\mbox{${\cal N}$}(\boldsymbol{\theta}|\mathbf{m}_{i},\mbox{$\mbox{$\mathbf{I}$}$})$. Then, $\mathbf{m}_{i}=\boldsymbol{\theta}_{i}$ if we estimate $q_{1}$ after using the delta method and products of all the Gaussian will simply give the addition as it is in Eq. 12. Therefore, TA can be seen as posterior correction with $q^{\text{iso}}$ but ignoring all the correction terms in Eq. 14. To improve TA, we can approximate the correction terms, for example, by using the $2^{\text{nd}}$-order Taylor expansion. This gives us a better Hessian-weighted version of TA. Essentially, we proceed similarly to Eq. 10 to get
@@
$\begin{split}\ell_{i}(\boldsymbol{\theta})-\hat{\ell}_{i|i}(\boldsymbol{\theta})&\approx\mbox{$\frac{1}{2}$}(\boldsymbol{\theta}-\mathbf{m}_{i})^{\top}\mathbf{H}_{i}(\boldsymbol{\theta}-\mathbf{m}_{i})\\ \ell_{\text{pre}}(\boldsymbol{\theta})-\hat{\ell}_{\text{pre}|\text{pre}}(\boldsymbol{\theta})&\approx\mbox{$\frac{1}{2}$}(\boldsymbol{\theta}-\mathbf{m}_{\text{pre}})^{\top}\mathbf{H}_{\text{pre}}(\boldsymbol{\theta}-\mathbf{m}_{\text{pre}})\end{split}$ (15)
@@
where $\mathbf{H}_{i}=\nabla^{2}\ell_{i}(\mathbf{m}_{i})$ and $\mathbf{H}_{\text{pre}}=\nabla^{2}\ell_{\text{pre}}(\mathbf{m}_{\text{pre}})$. Plugging these in Eq. 14, we get the following estimate of the merged mean, as shown in App. B,
@@
where $\mathbf{H}_{\text{ha}}=\mbox{$\mbox{$\mathbf{I}$}$}+\mathbf{H}_{\text{pre}}+\sum_{i}\alpha_{i}\mathbf{H}_{i}$.
This is exactly the update derived by Daheim et al. (2024, Eq. 12) by using gradient mismatch. In fact, it is easy to show that gradient mismatch is a special case of posterior correction. This is obtained by simply taking the derivatives of the correction term,
@@
Posterior correction generalizes their approach and relaxes the assumptions made by them (they assume a quadratic regularizer). Similarly to previous section, the result can also be directly obtained by expanding the posterior form from $q^{\text{iso}}$ to $q^{\text{full}}$.
@@ 3.4 Federated Learning as Posterior Correction
The goal of federated learning (McMahan et al., 2017) is to learn a joint model $\boldsymbol{\theta}_{\text{jnt}}$ at a server by communicating with local models $\boldsymbol{\theta}_{i}$ for $i=1,2,\ldots,t$ trained separately at $t$ clients. The local models use local losses $\ell_{i}$ which are not accessible at the server. This restriction forces the server to adapt all the available information and iteratively arrive at the correct solution. We will now show that the methods used in federated learning can be seen as an iterative posterior correction to eventually drive the error to zero. This is unlike the methods discussed so far which all used approximations for the correction term and therefore do not achieve perfect correction. Federated learning fixes this.
To show this, we start with a variational formulation for the joint-model learning and write the dual form of the solution using Eq. 3,
$\begin{split}q_{\text{jnt}}&=\arg\min_{q\in\mathcal{Q}}\,\,\mathbb{E}_{q}\left[\sum_{i=1}^{t}\ell_{i}\right]+\mathbb{D}_{\text{KL}}[q\,\|\,p_{0}]\\ &\,\,\propto p_{0}\prod_{i=1}^{t}\exp\left(-\hat{\ell}_{i|\text{jnt}}\right),\text{ where }\hat{\ell}_{i|\text{jnt}}(\boldsymbol{\theta})=\langle\mathbf{T}(\boldsymbol{\theta}),\widetilde{\nabla}\mathbb{E}_{q_{\text{jnt}}}[\ell_{i}]\rangle.\end{split}$ (18)
@@
Here, for simplicity, we have assume that $q_{\text{jnt}}$ and $p_{0}$ are both the same exponential family which implies $p_{0}\propto\exp(-\hat{\ell}_{0|\text{jnt}})$ and simplifies the discussion. However, the connection we discuss is valid even when this condition does not hold.
The dual form factorizes across local models which suggests a natural adaptation strategy to exploit the structure. Since each local ^$\hat{\ell}_{i|\text{jnt}}$ contributes a factor, we can learn a local model $q_{i}$ to estimate a local approximation ^$\hat{\ell}_{i|i}$ and slowly try to make it similar to the desired ^$\hat{\ell}_{i|\text{jnt}}$. To formalize this, consider the following estimate of $q_{\text{jnt}}$ obtained by using natural-gradients at local models $q_{i}$,
@@
$\hat{q}_{\text{jnt}}\propto p_{0}\prod_{i=1}^{t}\exp\left(-\hat{\ell}_{i|i}\right)\quad\text{ where }\hat{\ell}_{i|i}(\boldsymbol{\theta})=\langle\mathbf{T}(\boldsymbol{\theta}),\widetilde{\nabla}\mathbb{E}_{{\color[rgb]{1,0,0}q_{i}}}[\ell_{i}]\rangle.$ (19)
This differs from Eq. 18 because – instead of computing natural gradients at $q_{\text{jnt}}$ – here we compute them at a local $q_{i}$ (highlighted in red). We can then write the correction for the posterior $\hat{q}_{\text{jnt}}$ by plugging it in the first line of Eq. 18 and simplifying similarly to Eq. 4,
$q_{\text{jnt}}=\arg\min_{q\in\mathcal{Q}}\,\,\sum_{i=1}^{t}\mathbb{E}_{q}\left[\ell_{i}-\hat{\ell}_{i|i}\right]+\mathbb{D}_{\text{KL}}[q\,\|\,\hat{q}_{\text{jnt}}]\quad\implies q_{\text{jnt}}\propto\hat{q}_{\text{jnt}}\prod_{i=1}^{t}\exp\left(\ell_{i|\text{jnt}}-\hat{\ell}_{i|i}\right).$ (20)
A derivation for the second expression is included in App. C. The expression suggests that more accurate $\hat{q}_{\text{jnt}}$ yields smaller corrections, which is also faster to adapt to get to $q_{\text{jnt}}$. The goal of local $q_{i}$ should be to continue to reduce the correction. One straightforward strategy could be to locally minimize at each client $i$,
@@
$q_{i}=\arg\min_{q_{i}\in\mathcal{Q}}\,\,\mathbb{E}_{q_{i}}\left[\ell_{i}-\hat{\ell}_{i|i}\right]+\mathbb{D}_{\text{KL}}[q_{i}\,\|\,\hat{q}_{\text{jnt}}].$ (21)
We can then iterate between Eq. 19 and Eq. 21. This strategy is exactly the Partition Variational Inference (PVI) method of Ashman et al. (2022). For a comparison, see the version discussed in Swaroop et al. (2025, Eq. 5). Essentially, by using Eq. 19, we can see that each local update of Eq. 21 simply replaces the old factor ^$\hat{\ell}_{i|i}$ in $\hat{q}_{\text{jnt}}$ by a fresh one,

### v3

@@ 3.2.3 Memory Perturbation as Posterior Correction
$\boldsymbol{\lambda}_{t\backslash j}\approx\operatorname*{arg\,min}_{\boldsymbol{\mu}}\,\,-\langle\boldsymbol{\mu},\widetilde{\nabla}\mathbb{E}_{q_{t}}[\ell_{i}]\rangle+\mathbb{D}_{\text{KL}}[q\,\|\,q_{t}]\quad\implies\quad\boldsymbol{\lambda}_{t\backslash j}\approx\boldsymbol{\lambda}_{t}+\widetilde{\nabla}\mathbb{E}_{q_{t}}[\ell_{i}].$ (16)
@@
The second expression is obtained by simply taking the derivative with respect to $\boldsymbol{\mu}$ and setting it to zero. The derivation simplifies by noting that the derivative $\nabla_{\boldsymbol{\mu}}\mathbb{D}_{\text{KL}}[q_{t\backslash j}\,\|\,q_{t}]=\boldsymbol{\lambda}_{t\backslash j}-\boldsymbol{\lambda}_{t}$ (khanRue23, Eq. 23). The last expression above is precisely the Memory-Perturbation Equation shown in nickl2023memory. It is easy to check that when we use the $q^{\text{full}}$ family the equation above reduces to Eq. 15. The memory-perturbation equation gives a general way to expand the class of the variational family to get better estimators, and as shown above it is a special instance of posterior correction where corrections are simply ignored. Overall, these connections suggests two ways to improve influence estimation: either we can expand the posterior-family class, or we could explicitly reduce the correction.
@@ 3.3 Model Merging as Posterior Correction
Model merging is a popular technique to improve capabilities of LLMs without retraining them from scratch. Given a pre-trained model with parameters $\boldsymbol{\theta}_{\text{pre}}$ on loss $\ell_{\text{pre}}$, we ‘fine-tune’ models $\boldsymbol{\theta}_{i}$ on individual tasks $\ell_{i}$. For instance, given an LLM trained on English text, we may fine-tune it on other languages and merge them all to get a model that can also generate texts in other languages. The fine-tuning procedure is essentially identical to the standard stochastic training but is initialized at $\boldsymbol{\theta}_{\text{pre}}$ and trained only on $\ell_{i}$ with early stopping to avoid overfitting. Given fine-tuned models $\boldsymbol{\theta}_{i}$ for all tasks $i$, the merging is done via a simple addition, for instance, Task-Arithmetic (TA) (IlharcoRWSHF23) uses the following merging (with weights $\alpha_{i}>0$):
$\boldsymbol{\theta}_{\text{ta}}=\boldsymbol{\theta}_{\text{pre}}+\sum_{i=1}^{t}\alpha_{i}(\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{\text{pre}}).$ (17)
This simple merging works extremely well for unknown reasons. In what follows, we will derive it as a special case of posterior correction where correction is ignored, and then show that an explicit reduction of the correction leads to a better merging method.
3.3.1 Posterior Merging via Bayesian Arithmetic
Similarly to other sections, we start by formulating model merging as a variational problem. We denoteby $q_{\text{pre}}$ the pre-trained posterior on $\ell_{\text{pre}}$ and by $q_{i}$ the posterior obtained by fine-tuning $q_{\text{pre}}$ over the task $\ell_{i}$. For fine-tuning, we use the following variational objective as a proxy where the KL term is used to mimic the initialization used to warm start,
$q_{i}=\operatorname*{arg\,min}_{q\in\mathcal{Q}^{\text{iso}}}\,\,\mathbb{E}_{q}[\ell_{i}]+\mathbb{D}_{\text{KL}}[q\,\|\,q_{\text{pre}}].$ (18)
Given $q_{\text{pre}}$ and all $q_{i}$, our goal is to obtain the model trained from scratch on all the data jointly. A simple rule to merge the posteriors $q_{i}$ is the following:
@@ 3.3.1 Posterior Merging via Bayesian Arithmetic
which has a similar form to Eq. 17 if we replace addition and subtraction by multiplication and division. Due to this analogy, we call this posterior merging as ‘Bayesian Arithmetic (BA)’. The above rule is commonly used in various distributed Bayesian computation, such as, Bayesian data fusion (mutambara2019decentralized; durrant2001data; wu2022bayesian), Bayesian Committee Machine (tresp2000bayesian), Consensus Monte Carlo (scott2022bayes), and approximate inference (vehtari2020expectation; ashman2022partitioned).
3.3.2 Merging as Posterior Correction
Clearly, posterior merging does not recover the joint posterior over the whole data. We will now use posterior correction to write an exact expression for the error. Because both $q_{i}$ and $q_{\text{pre}}$ are obtained by training a variational objetive, we can write their dual from as follows:
$q_{i}\propto q_{\text{pre}}\exp(-\hat{\ell}_{i})\qquad\text{and}\qquad q_{\text{pre}}\propto\exp(-\hat{\ell}_{\text{pre}}),$
where $\hat{\ell}_{i}$ is the site for $\ell_{i}$ evaluated at $q_{i}$ and $\hat{\ell}_{\text{pre}}$ is the same for the whole pre-trained loss $\ell_{\text{pre}}$ (we assume that it also contains the prior used during pre-training). The more corret notation would be to use $\hat{\ell}_{i|i}$, but we use $\hat{\ell}_{i}$ which is simpler. Using a procedure similar to Eq. 4, we can rewrite the joint posterior shown in the second line,
@@ 3.3.2 Merging as Posterior Correction
$\begin{split}q_{\text{retrained}}&=\operatorname*{arg\,min}_{q\in\mathcal{Q}}\,\,\mathbb{E}_{q}\left[\ell_{\text{pre}}+\sum_{i=1}^{t}\alpha_{i}\ell_{i}\right]+\mathbb{D}_{\text{KL}}[q\,\|\,p_{0}]\\ &=\operatorname*{arg\,min}_{q\in\mathcal{Q}}\,\,\mathbb{D}_{\text{KL}}[q\,\|\,q_{\text{ba}}]+\sum_{i=1}^{t}\alpha_{i}\underbrace{\mathbb{E}_{q}[\ell_{i}-\hat{\ell}_{i}]}_{\text{Task correction}}+\underbrace{\mathbb{E}_{q}[\ell_{\text{pre}}-\hat{\ell}_{\text{pre}}]}_{\text{LLM correction}}.\end{split}$ (20)
The result is obtained by simply multiplying and dividing by $q_{\text{ba}}$ as we did in Eq. 4. The last two terms contain two kinds of corrections: the first one contains corrections to tackle the interference due to fine-tuned posteriors, while the second correction is needed to mitigate the interference to the original LLM. The reformulation above suggest to addressed both types of corrections to effectively merge the posteriors.
3.3.3 Task Arithmetic and Hessian-Based Merging as Posterior Correction
It is easy to check that the TA merging in Eq. 17 can be recovered as a special case if we make two changes. First, we restrict the set of distributions to the $q^{\text{iso}}$ family, setting $q_{\text{pre}}=\mbox{${\cal N}$}(\boldsymbol{\theta}|\mathbf{m}_{\text{pre}},\mbox{$\mbox{$\mathbf{I}$}$})$ and all $q_{i}=\mbox{${\cal N}$}(\boldsymbol{\theta}|\mathbf{m}_{i},\mbox{$\mbox{$\mathbf{I}$}$})$ for $i=1,2,\ldots,t$. Second, we ignore the correction term. With these approximations, solving the above gives us $q_{\text{retrained}}=q_{\text{ba}}$. The solution is then simply obtained by using Eq. 19 which yields $\mathbf{m}_{\text{ba}}=\mathbf{m}_{\text{pre}}+\sum_{i=1}^{t}\alpha_{i}(\mathbf{m}_{i}-\mathbf{m}_{\text{pre}})$, a merging equivalent to Eq. 17 if we use the delta method to train both the LLM and perform fine-tuning. By ignoring the correction term, we obtain task arithmetic as a special instance of Bayesian arithmetic on the $q^{\text{iso}}$ family.
To get a better merging strategy, we can now explicitly approximate the correction terms, instead of ignoring it. We will use the same strategy as Sec. 3.2.2 relying on the second-order Taylor expansion to approximate both the task and LLM corrections,
@@ 3.3.3 Task Arithmetic and Hessian-Based Merging as Posterior Correction
$\begin{split}\ell_{i}(\boldsymbol{\theta})-\hat{\ell}_{i}(\boldsymbol{\theta})&\approx\mbox{$\frac{1}{2}$}(\boldsymbol{\theta}-\mathbf{m}_{i})^{\top}\mathbf{H}_{i}(\boldsymbol{\theta}-\mathbf{m}_{i})\\ \ell_{\text{pre}}(\boldsymbol{\theta})-\hat{\ell}_{\text{pre}}(\boldsymbol{\theta})&\approx\mbox{$\frac{1}{2}$}(\boldsymbol{\theta}-\mathbf{m}_{\text{pre}})^{\top}\mathbf{H}_{\text{pre}}(\boldsymbol{\theta}-\mathbf{m}_{\text{pre}})\end{split}$ (21)
@@
where $\mathbf{H}_{i}=\nabla^{2}\ell_{i}(\mathbf{m}_{i})$ and $\mathbf{H}_{\text{pre}}=\nabla^{2}\ell_{\text{pre}}(\mathbf{m}_{\text{pre}})$. Plugging these in Eq. 20, we get the following estimate of the merged mean, as shown in App. F,
@@
where $\mathbf{H}_{\text{ha}}=\mbox{$\mbox{$\mathbf{I}$}$}+\mathbf{H}_{\text{pre}}+\sum_{i}\alpha_{i}\mathbf{H}_{i}$. This is exactly the update derived by DaheimMPGK24 by using gradient mismatch. In fact, it is easy to show that gradient mismatch is a special case of posterior correction. This is obtained by simply taking the derivatives of the correction term with respect to the mean $\mathbf{m}$,
@@
The second equality follows by using Bonnet’s theorem $\nabla_{\mathbf{m}}\mathbb{E}_{q}[\ell_{i}]=\mathbb{E}_{q}[\nabla\ell_{i}]$ and the approximation afterward is due to the delta method. The correction term in this case account for gradient mismatches which can be seen as a special case of the natural-gradient mismatch used in this paper. Posterior correction generalizes their approach and relaxes the assumptions made by them (they assume a quadratic regularizer and no such restrictions are needed in our derivation).
An alternate way to derive the same result is to expand the size of the family from $q^{\text{iso}}$ to $q^{\text{full}}$, as we did in the previous section. Similar to the previous derivation, if we ignore the corrections, then merging reduces to the computation of $q_{\text{ba}}$. Then, by denoting $q_{i}=\mbox{${\cal N}$}(\boldsymbol{\theta}|\mathbf{m}_{i},\mbox{$\mbox{$\mathbf{S}$}$}_{i}^{-1})$ and $q_{\text{pre}}=\mbox{${\cal N}$}(\boldsymbol{\theta}|\mathbf{m}_{\text{pre}},\mbox{$\mbox{$\mathbf{S}$}$}_{\text{pre}}^{-1})$, we get the following merging formula,
$\mathbf{m}_{\text{ba}}=\mathbf{m}_{\text{pre}}+\sum_{i=1}^{t}\alpha_{i}\mbox{$\mbox{$\mathbf{S}$}$}_{\text{ba}}^{-1}(\mbox{$\mbox{$\mathbf{I}$}$}+\mbox{$\mbox{$\mathbf{S}$}$}_{i})(\mathbf{m}_{i}-\mathbf{m}_{\text{pre}}),$ (24)
where $\mbox{$\mbox{$\mathbf{S}$}$}_{\text{ba}}=\mbox{$\mbox{$\mathbf{I}$}$}+\mbox{$\mbox{$\mathbf{S}$}$}_{\text{pre}}+\sum_{i}\alpha_{i}\mbox{$\mbox{$\mathbf{S}$}$}_{i}$. Essentially, the Hessians in Eq. 22 are now replaced by the precision matrices. If we use the delta method to optimize over the $q^{\text{full}}$ family, then this will reduce to Hessian-based merging. This again shows that we do not have to explicitly approximate the correction and a reduction naturally occurs when we use a more flexible posterior form, yielding a better adaptation method automatically.
@@ 3.4 Federated Learning as Posterior Correction
The goal of distributed and federated learning (mcmahan2016fedavg) is to learn a joint model $\boldsymbol{\theta}_{\text{jnt}}$ at a server by communicating with local models $\boldsymbol{\theta}_{i}$ for $i=1,2,\ldots,t$ trained separately at $t$ clients. The local models use local losses $\ell_{i}$ not accessible at the server. This restriction forces the server to adapt all the available information and iteratively arrive at the correct solution. We will now show that the methods used in federated learning can be seen as an iterative posterior-correction to eventually drive the corrections to zero. This is unlike the methods discussed so far which all used approximations for the correction term and therefore do not achieve perfect correction. Algorithms discussed here fix this issue.
3.4.1 Federated Posterior Correction
@@
To show this, we start with a variational formulation for the joint-model learning. For simplicity, we will assume that $q_{\text{jnt}}$ and $p_{0}$ belong to the same exponential family because it simplifies the calculation, but the discussion is valid even when this condition does not hold. The variational formulation is shown below,
@@
$\begin{split}q_{\text{jnt}}^{*}&=\operatorname*{arg\,min}_{q\in\mathcal{Q}}\,\,\mathbb{E}_{q}\left[\sum_{i=1}^{t}\ell_{i}\right]+\mathbb{D}_{\text{KL}}[q\,\|\,p_{0}]\\ &\,\,\propto p_{0}\prod_{i=1}^{t}\exp\left(-\hat{\ell}_{i|\text{jnt}}^{*}\right),\text{ where }\hat{\ell}_{i|\text{jnt}}^{*}(\boldsymbol{\theta})=\langle\mathbf{T}(\boldsymbol{\theta}),\widetilde{\nabla}\mathbb{E}_{q_{\text{jnt}}^{*}}[\ell_{i}]\rangle.\end{split}$ (25)
In the second line, we write the solution in its dual form, similarly to Eq. 3. The form factorizes across local models which suggests a natural adaptation strategy by exploiting the structure. Since each local $\hat{\ell}_{i|\text{jnt}}^{*}$ contributes a factor, we can slowly try to move the locally learned approximations towards the desired $\hat{\ell}_{i|\text{jnt}}^{*}$. As we show, federated learning can be interpreted as following this recipe.
To show this, we first present a federated algorithm for posterior correction from which we can derive other federated learning algorithms. We denote by $\hat{\ell}_{i}$ a local site of $\ell_{i}$ at the local $q_{i}$. The algorithm starts with an initial value of $\hat{\ell}_{i}=\langle\mathbf{T}(\boldsymbol{\theta}),\hat{\boldsymbol{\lambda}}_{i}\rangle$ with a first guess for $\hat{\boldsymbol{\lambda}}_{i}$. Then the algorithm follows a double loop structure. In the outer loop, we construct a joint posterior as follows,
$q_{\text{jnt}}\leftarrow\frac{1}{\mathcal{Z}_{\text{jnt}}}p_{0}\prod_{i=1}^{t}\exp\left(-\hat{\ell}_{i}\right).$ (26)
Then, we use it to write a posterior-corrected version of Eq. 25 (similarly to Eq. 4),
$q_{\text{jnt}}^{*}=\operatorname*{arg\,min}_{q\in\mathcal{Q}}\,\,\sum_{i=1}^{t}\mathbb{E}_{q}\left[\ell_{i}-\hat{\ell}_{i}\right]+\mathbb{D}_{\text{KL}}[q\,\|\,q_{\text{jnt}}].$ (27)
In the inner loop, we separately optimize with respect to each $q_{i}$ by solving
@@
$q_{i}\leftarrow\operatorname*{arg\,min}_{q_{i}\in\mathcal{Q}}\,\,\frac{1}{\rho}\mathbb{E}_{q_{i}}\left[\ell_{i}-\hat{\ell}_{i}\right]+\mathbb{D}_{\text{KL}}[q_{i}\,\|\,q_{\text{jnt}}].$ (28)
Here, we multiply the local correction by $\rho$ as a simple approximation of the whole sum; for instance, we can set $\rho=1/t$ which is equivalent to weighting the sum by $t$. The site parameter is updated to be the natural gradients at the new $q_{i}$, that is, $\hat{\boldsymbol{\lambda}}_{i}\leftarrow\widetilde{\nabla}\mathbb{E}_{q_{i}}[\ell_{i}]$. The outer and inner loops in Eq. 26 and Eq. 28 respectively are repeated until convergence. We call this algorithm the Federated Posterior Correction algorithm. It is easy to check that the solution in Eq. 25 is recovered upon convergence; see Sec. G.1. As a result, upon convergence, the correction term also vanishes and $\hat{\ell}_{i}$ converge to $\hat{\ell}_{i|\text{jnt}}^{*}$.
3.4.2 Alternating Minimization Algorithm as Posterior Correction
We will now show how existing federated learning algorithms can be seen as an implementation of the above federated posterior-correction algorithm. The closest algorithm to ours is perhaps the Alternating Minimization Algorithm (AMA) of Ts91 which uses splitting methods for optimization. For instance, consider the following objective
$\boldsymbol{\theta}_{\text{jnt}}^{*}=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\sum_{i=1}^{t}\ell_{i}(\boldsymbol{\theta})+\mbox{$\frac{1}{2}$}\|\boldsymbol{\theta}\|^{2}.$ (29)
A quadratic regularizer is used above for simplicity but the method can be used for a much larger class of objectives.
A detailed description of AMA is given in Sec. G.2. Below, we summarize the updates and then compare them to our federated algorithm. In AMA, we define local loss functions $\ell_{i}(\boldsymbol{\theta}_{i})$ over local parameters $\boldsymbol{\theta}_{i}$ for all $i=1,2,\ldots,t$. A global parameter $\boldsymbol{\theta}_{\text{jnt}}$ is also introduced which operates over the quadratic regularizer $\|\boldsymbol{\theta}_{\text{jnt}}\|^{2}/2$. Then a Lagrangian is formed to enforce constraints $\boldsymbol{\theta}_{i}=\boldsymbol{\theta}_{\text{jnt}}$ for all $i$ via dual vectors $\mbox{$\mbox{$\mathbf{v}$}$}_{i}$ (also known as Lagrange multipliers). The Lagrangian gives rise to the following update for the triplet $(\boldsymbol{\theta}_{\text{jnt}},\boldsymbol{\theta}_{i},\mbox{$\mbox{$\mathbf{v}$}$}_{i})$,
$\boldsymbol{\theta}_{\text{jnt}}\leftarrow-\sum_{i=1}^{t}\mbox{$\mbox{$\mathbf{v}$}$}_{i},\quad\quad\boldsymbol{\theta}_{i}\leftarrow\operatorname*{arg\,min}_{\boldsymbol{\theta}_{i}}\,\,\ell_{i}(\boldsymbol{\theta}_{i})-\boldsymbol{\theta}_{i}^{\top}\mbox{$\mbox{$\mathbf{v}$}$}_{i}+\rho\mbox{$\frac{1}{2}$}\|\boldsymbol{\theta}_{i}-\hat{\boldsymbol{\theta}}_{\text{jnt}}\|^{2},\quad\quad\mbox{$\mbox{$\mathbf{v}$}$}_{i}\leftarrow\nabla\ell_{i}(\boldsymbol{\theta}_{i}).$ (30)
To show the similarity to our federated algorithm, we rewrite Eqs. 26 and 28 in a similar form where we update the triplet $(\boldsymbol{\lambda}_{\text{jnt}},\boldsymbol{\mu}_{i},\hat{\boldsymbol{\lambda}}_{i})$,
$\boldsymbol{\lambda}_{\text{jnt}}\leftarrow-\sum_{i=0}^{t}\hat{\boldsymbol{\lambda}}_{i},\quad\quad\boldsymbol{\mu}_{i}\leftarrow\operatorname*{arg\,min}_{\boldsymbol{\mu}_{i}}\mathbb{E}_{q_{i}}[\ell_{i}]-\langle\boldsymbol{\mu}_{i},\hat{\boldsymbol{\lambda}}_{i}\rangle+\rho\mathbb{D}_{\text{KL}}[q_{i}\,\|\,q_{\text{jnt}}],\quad\hat{\boldsymbol{\lambda}}_{i}\leftarrow\widetilde{\nabla}\mathbb{E}_{q_{i}}[\ell_{i}].$ (31)
A derivation is in Sec. G.3. Essentially, the global natural parameter $\boldsymbol{\lambda}_{\text{jnt}}$ is updated by summing the local site parameters $\hat{\boldsymbol{\lambda}}_{i}$, which are the natural gradients of the expected local-loss functions. The role of the dual vector $\mbox{$\mbox{$\mathbf{v}$}$}_{i}$ is now played by the site parameters $\hat{\boldsymbol{\lambda}}_{i}$ (a minor detail is that $\hat{\boldsymbol{\lambda}}_{0}$ is the natural parameter of $p_{0}$). The second update is performed in the $\boldsymbol{\mu}_{i}$-space where the linear term $\boldsymbol{\theta}^{\top}\mbox{$\mbox{$\mathbf{v}$}$}_{i}$ is now replaced by the $\langle\boldsymbol{\mu}_{i},\hat{\boldsymbol{\lambda}}_{i}\rangle$. The theorem below shows AMA can be recovered as a special case.
Theorem 1
For $q^{\text{iso}}$ family, Eq. 31 reduces to Eq. 30 if we approximate $\mathbb{E}_{q_{i}}[\ell_{i}]\approx\ell_{i}(\mathbf{m}_{i})$.
A detailed proof is given in Sec. G.3. Overall, these results show that not only can we recover AMA from our algorithm, we can also generalize it by choosing more flexible exponential-family forms.
3.4.3 Partition Variational Inference as Posterior Correction
Our algorithms closely resembles Partitioned Variational Inference (PVI) (ashman2022partitioned). This also highlights some potential issues with the PVI algorithm and provides ways to fix them. PVI aims for distributed federated Bayesian inference where the losses $\ell_{i}$ are simply negative of log-likelihoods, that is, $\log\text{lik}_{i}=-\ell_{i}$. PVI follows the strategy of the well-known Expectation Propagation (EP) algorithm, where local site functions, denoted by $s_{i}$, are used to approximate the likelihood, that is, $s_{i}\approx\text{lik}_{i}$. A major difference is that the PVI algorithm does not assume that $q$ is an EF posterior, so $s_{i}$ may not necessarily be defined in terms of natural gradients. In such cases, PVI implements the following steps, starting with $q_{\text{jnt}}\leftarrow p_{0}$ and $s_{i}\leftarrow 1$,